OpenStudy (anonymous):

solve the equation. check for extraneous solutions 4|4-3x|=4x+6

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

|x| = a implies x = +a or, x = -a

OpenStudy (anonymous):

i only got the first answer which is 11/8 but there is a second answer and i cant get it

OpenStudy (anonymous):

cud u show me how did u get the first one?

OpenStudy (anonymous):

4|4-3x|=4x+6 16-12x=4x+6 +12x with 4x +16 with the 6 16x/16=22/16 x=11/8

OpenStudy (anonymous):

nah u did it wrong

OpenStudy (anonymous):

now i cant get the second answer

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

wait then how did u do it

OpenStudy (anonymous):

16-12x=4x+6 12x + 4x = 16 - 6 16x = 10 x = 10/16 = 5/8

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

yes so then how would you get the solution for the second one

OpenStudy (anonymous):

the only choices with 5/8 is x=5/8 or x= 11/8 x=5/8 or x=11/4

OpenStudy (anonymous):

the other one has a negative sign so i figured that's not it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and

OpenStudy (anonymous):

if u want to kno all the choices i can tell u

OpenStudy (anonymous):

cant u use the same method u used for the other one with this one? it's the same process right?

OpenStudy (anonymous):

or since it's a different part of the problem would u have to use a different method?

OpenStudy (anonymous):

helloooo?

OpenStudy (anonymous):

forget it i got the answer already but thnx